OpenStudy (zarvox1):

Find the perimeter and area of a rectangle with a length of 3.2 m and a width of 1.5 m.

OpenStudy (anonymous):

The perimeter of a rectangle is p = 2l + 2 w. So, 3.2+3.2+1.5+1.5= 9.4m The area of a rectangle is a =lw. So, 3.2 x 1.5 = 4.8m^2

OpenStudy (anonymous):

9.4 m \[3.2 \times 2 = 6.4 + 1.5 \times 2 =3 so 6.4 +3 =9.4\]